### v1

# Weil Restriction and the Motivic Cycle Class Map

Abstract.
The étale cycle class map links Chow groups with étale cohomology. In the framework of motivic cohomology, we have the motivic cycle class map, which appears as a comparison morphism relating motivic and Lichtenbaum cohomology. By an examination of the comparison morphism constructed by Geisser and Levine, we show that the motivic cycle class map is induced by the étale cycle class map. We then study the behavior of cycle class maps under Weil restriction. Extending Karpenko’s construction of the Weil restriction for algebraic cycles and Chow groups, we introduce a corresponding Weil restriction for $\ell$-adic cohomology and prove its compatibility with the motivic cycle class map. This yields a conceptual explanation of the descent of cycle classes under finite Galois extensions.
Key words and phrases:
Weil restriction; Motivic cohomology; Étale realization; Chow rings. Mathematical Subject Classification (2020): 14F42; 14C25
Cohomology theories of algebraic schemes occupy a central position in modern algebraic geometry. A remarkable example is Deligne’s proof of the Weil conjecture, in which he made essential use of deep properties of étale cohomology.
This naturally leads to the question of how different cohomology theories are related. As a bridge between such theories, the étale cycle class map provides an initial link between Chow theory—a Borel–Moore oriented cohomology theory in the sense of Levine and Morel [9]—and étale cohomology.
More concretely, for a smooth variety $X$ of finite type over a field $k$, there is a ring homomorphism
$cl_{X}^{*}\colon\mathop{\mathrm{CH}}\nolimits^{*}(X)\longrightarrow H^{2*}\!\left(X_{\acute{e}t},\mathbb{Q}_{\ell}(*)\right),$
under which the intersection product on $\mathop{\mathrm{CH}}\nolimits^{*}(X)$ corresponds to the cup product in étale cohomology. Here $\mathbb{Q}_{\ell}(*)$ denotes the $\ell$-adic coefficient system equipped with the appropriate Tate twist.
Another important construction in algebraic geometry is the Weil restriction, a powerful tool for descent with numerous applications in arithmetic and geometry. The Weil restriction for algebraic cycles—and for Chow groups—was established by Karpenko [7]. In parallel with Karpenko’s work, we construct the Weil restriction map for $\ell$-adic cohomology.
The relationship between Chow groups and étale cohomology can be further clarified within the framework of motivic cohomology—conjectured by Beilinson and Lichtenbaum and later constructed by Voevodsky. In this setting, the Chow theory and étale cohomology are connected naturally as a comparison morphism arising from the change of topology from the Nisnevich site to the étale site, relating motivic cohomology to Lichtenbaum cohomology, which is known as the motivic cycle class map.
In this paper, we show that the the motivic cycle class map is indeed induced by the étale cycle class map. This is achieved by a careful examination of the construction of the comparison morphism due to Geisser and Levine [3] using Bloch’s cycle complex.
We then investigate the compatibility between Chow groups and étale cohomology under Weil restriction. In particular, we study this compatibility within the framework of motivic cohomology.
Notation and conventions
Throughout this paper, schemes are assumed to be separated and of finite type over a field. A variety means an integral scheme, i.e., an irreducible and reduced scheme.
For an abelian group $\Lambda$, we write $\underline{\Lambda}$ for the associated constant étale sheaf.
For a scheme $X$, $H^{*}(X,\mathcal{F})$ denotes the cohomology groups of a sheaf $\mathcal{F}$ (on the relevant site), and for a closed subscheme $Z\subset X$, $H^{*}_{Z}(X,\mathcal{F})$ denotes the cohomology groups with support in $Z$.
For two sheaves $\mathcal{F}$ and $\mathcal{G}$ on a given site, we write $\operatorname{Hom}(\mathcal{F},\mathcal{G})$ for the group of global morphisms and $\mathcal{H}om(\mathcal{F},\mathcal{G})$ for the internal Hom sheaf.
1. Preliminary
1.1. Weil restriction of schemes
We briefly review basics of Weil restriction of schemes. A detailed discussion can be found in [6] or [1].
Let $L/k$ be a finite Galois field extension, and let $X$ be a scheme over $L$. The Weil restriction of $X$ along the extension $L/k$, denoted by $\mathcal{R}_{L/k}(X)$ (or simply $\mathcal{R}(X)$ when no confusion arises), is a $k$-scheme (if exists) characterized by the following universal property
$\operatorname{Hom}_{k}\bigl(S,\mathcal{R}(X)\bigr)\;\cong\;\operatorname{Hom}_{L}\bigl(S\times_{k}L,\,X\bigr),$
for any $k$-scheme $S$. Weil restriction always exists when considering quasi-projective schemes.
Under the same assumptions as above, we may describe $\mathcal{R}(X)_{L}$ more explicitly. Let $G:=\mathop{\mathrm{Gal}}(L/k)$ be the Galois group corresponding to the field extension. For any element $\sigma\in G$, we denote by $X^{\sigma}$ the conjugate scheme of $X$, which is obtained by composing the structure morphism of $X$ with $\sigma:\operatorname{Spec}(L)\to\operatorname{Spec}(L)$, the field automorphism induced by $\sigma$. Then there is a canonical isomorphism
(1.1) $\mathcal{R}(X)_{L}\;\cong\;\prod_{\sigma\in G}X^{\sigma}.$
Weil restriction enjoys several functorial properties, we list some of them in the following.
Proposition 1.2 ([7, Proposition 1.1]).
Let $L/k$ be a finite Galois extension, and assume that the Weil restriction of schemes exists in the following setting. Then:
  • (1)
    For schemes $X,Y$ over $L$, we have
    $\mathcal{R}(X\times_{L}Y)\cong\mathcal{R}(X)\times_{k}\mathcal{R}(Y).$
  • (2)
    For the closed embedding of schemes over $L$, $i\colon X\hookrightarrow Y$, the induced morphism
    $\mathcal{R}(i)\colon\mathcal{R}(X)\to\mathcal{R}(Y)$
    is also a closed embedding.
  • (3)
    If $X$ is a smooth scheme over $L$, then $\mathcal{R}(X)$ is smooth over $k$.
1.2. Weil restriction of algebraic cycles
Now we may turn to Weil restriction of algebraic cycles. The original construction is given by Karpenko in [7]. We briefly recall the construction for later use.
Recall that the group of algebraic cycles on a scheme $X$ is defined as the free abelian group generated by all closed subvarieties of $X$. We denote the group by $Z^{*}(X)$ or simply $Z(X)$. This group carries a natural grading by codimension, and we therefore write
$Z^{*}(X)=\bigoplus_{r=0}^{\dim X}Z^{r}(X),$
where $Z^{r}(X)$ denotes the subgroup generated by integral closed subschemes of codimension $r$ in $X$. Moreover, by quotienting out by the relation of rational equivalence (see [2]), we obtain the (integral) Chow group of $X$, denoted by $\mathop{\mathrm{CH}}\nolimits^{*}(X)$ (or simply $\mathop{\mathrm{CH}}\nolimits(X)$). We write $\mathop{\mathrm{CH}}\nolimits^{r}(X)$ for the $r$th graded piece of $\mathop{\mathrm{CH}}\nolimits^{*}(X)$, which is the quotient of $Z^{r}(X)$ by rational equivalence. We describe functorial properties of algebraic cycles with respect to certain morphisms.
For a flat morphism $f\colon Y\to X$ of constant relative dimension $d$, there is a pullback homomorphism
$f^{*}\colon Z^{r}(X)\to Z^{r}(Y),$
and hence $f^{*}\colon Z^{*}(X)\to Z^{*}(Y)$.
For a proper morphism $g\colon X\to Y$, there is a pushforward homomorphism
$g_{*}\colon Z^{r}(X)\to Z^{r-(\dim X-\dim Y)}(Y),$
and thus $g_{*}\colon Z^{*}(X)\to Z^{*-(\dim X-\dim Y)}(Y)$.
Pullbacks and pushforwards are functorial with respect to composition. Moreover, they respect the relation of rational equivalence; therefore, both operations descend to homomorphisms on Chow groups (see [2, Chapter I]).
Now we can start the construction of Weil restriction of algebraic cycles. Let $L/k$ be a finite Galois extension and let $X$ be a smooth quasi-projective scheme over $L$. The Weil restriction $\mathcal{R}(X)$ exists and is a smooth quasi-projective scheme over $k$ [7, Corollary 1.4]. Let $G=\mathop{\mathrm{Gal}}(L/k)$ be the Galois group.
Consider the natural projection morphism
$p\colon\mathcal{R}(X)_{L}\to\mathcal{R}(X).$
It induces an injective pullback homomorphism
$p^{*}\colon Z(\mathcal{R}(X))\longrightarrow Z(\mathcal{R}(X)_{L}).$
The key point is that the image of $p^{*}$ is precisely the subgroup of $G$-invariant cycles in $Z(\mathcal{R}(X)_{L})$ (see [7, Lemma 2.1]). In other words, there is an isomorphism
$p^{*}\bigl(Z(\mathcal{R}(X))\bigr)\cong Z(\mathcal{R}(X)_{L})^{G}.$
For any algebraic cycle $\alpha\in Z^{r}(X)$ and any $\sigma\in G$, we denote by $\alpha^{\sigma}$ the cycle on $X^{\sigma}$ obtained by base change via $\sigma$. Then
$\prod_{\sigma\in G}\alpha^{\sigma}$
defines a cycle in $Z\bigl(\mathcal{R}(X)_{L}\bigr)$ via the exterior product. This cycle is $G$-invariant, and therefore corresponds to a unique cycle $\mathcal{R}(\alpha)\in Z^{dr}(\mathcal{R}(X))$ such that
$p^{*}\bigl(\mathcal{R}(\alpha)\bigr)=\prod_{\sigma\in G}\alpha^{\sigma}.$
As a result, we obtain a well-defined map
$\mathcal{R}\colon Z^{r}(X)\longrightarrow Z^{dr}\bigl(\mathcal{R}(X)\bigr),$
where $d=[L:k]$. By taking the direct sum over all graded components, we obtain a map (Not a group homomorphism)
$\mathcal{R}\colon Z^{*}(X)\longrightarrow Z^{d\cdot*}(\mathcal{R}(X)),$
which we call the Weil restriction of algebraic cycles. The Weil restriction map is compatible with flat pullbacks, proper pushforwards, and both interior and exterior products of cycles [7, Proposition 2.5].
Moreover, the Weil restriction of algebraic cycles is compatible with the relation of rational equivalence [7, Corollary 3.3]. Hence, the map descends to the level of Chow groups,
$\mathcal{R}\colon\mathop{\mathrm{CH}}\nolimits(X)\longrightarrow\mathop{\mathrm{CH}}\nolimits(\mathcal{R}(X)).$
It remains compatible with pushforwards, pullbacks, products of cycles, and Gysin pullbacks on Chow groups. Finally, it is compatible with the intersection product [7, Proposition 3.4]. For a modern proof of compatibility with the intersection product, we refer the reader to [8], where Fulton’s deformation to the normal cone is used.
1.3. Étale cohomology and cycle class map
In this section, we recall some basics on étale cohomology and the construction of the étale cycle class map. This material is covered in standard expositions on étale cohomology; for example, Milne’s Étale Cohomology [11], which treats primarily the case over an algebraically closed field. In fact, the main statements remain valid over an arbitrary base field. For a more general and comprehensive treatment, we refer the reader to Cycle, SGA $4\frac{1}{2}$ [5].
We fix a base field $k$. Throughout this section, a sheaf is an étale sheaf; any cohomology group will mean étale cohomology, defined as sheaf cohomology on the étale site unless otherwise specified. Recall that for an abelian group $A$, we denote by $\underline{A}$ the associated constant étale sheaf.
Recall the following definition of Tate twist.
Definition 1.3.
[11, §2, Chapter V]
[…]
$\mathcal{F}(r):=\begin{cases}\mathcal{F}\otimes\mu_{n}^{\otimes r}&\text{if }r>0,\\ \mathcal{F}&\text{if }r=0,\\ \mathcal{H}om(\mathcal{F}(-r),\,\mathbf{\mathbb{Z}}/l\mathbb{Z})&\text{if }r<0.\end{cases}$
Now we turn to the construction of the étale cycle class map.
Let $X$ be a smooth variety over $k$, and let $\Lambda:=\mathbb{Z}/\ell\mathbb{Z}$ with $\ell\neq\mathrm{char}(k)$. For a smooth closed subvariety $i\colon Z\hookrightarrow X$ of codimension $r$, we have the Gysin map (see [11, §5, Chapter VI])
$i_{*}\colon H^{0}\!\left(Z,i^{*}\underline{\Lambda}\right)\longrightarrow H^{2r}_{Z}\!\left(X,\underline{\Lambda}(r)\right).$
Since $H^{0}(Z,i^{*}\underline{\Lambda})\cong\mathbb{Z}/\ell\mathbb{Z}$, the image of $1\in H^{0}(Z,i^{*}\underline{\Lambda})$ defines a class in $H^{2r}_{Z}(X,\underline{\Lambda}(r))$. Composing with the canonical map
$H^{2r}_{Z}(X,\underline{\Lambda}(r))\longrightarrow H^{2r}(X,\underline{\Lambda}(r)),$
we obtain a cohomology class in $H^{2r}(X,\underline{\Lambda}(r))$, which we denote by $cl_{X}^{r}(Z)$.
For a possibly singular closed subvariety $Z^{\prime}\subset X$, one can still define such a class using semi-purity [5, Proposition 2.2.6], which yields an identification
$H^{2r}_{Z^{\prime}}(X,\underline{\Lambda}(r))\;\cong\;H^{2r}_{Z^{\prime}\setminus Sing(Z^{\prime})}\!\bigl(X\setminus Sing(Z^{\prime}),\underline{\Lambda}(r)\bigr),$
where $Sing(Z^{\prime})$ denotes the singular locus of $Z^{\prime}$. This reduces the construction to the case of a closed embedding of smooth schemes
$i^{\prime}\colon Z^{\prime}\setminus Sing(Z^{\prime})\hookrightarrow X\setminus Sing(Z^{\prime}).$
Finally, by extending the construction linearly, we obtain the $r$th étale cycle class map
(1.4) $cl_{X}^{r}\colon Z^{r}(X)\longrightarrow H^{2r}\!\left(X,\underline{\Lambda}(r)\right).$
By taking the direct sum over all codimensions $r\geq 0$, we obtain the total étale cycle class map
(1.5) $cl_{X}\colon Z^{*}(X)\longrightarrow H^{2*}(X):=\oplus_{i=0}^{dim(X)}H^{2i}\!\left(X,\underline{\Lambda}(i)\right).$
The étale cycle class map enjoys good compatibility with operations on algebraic cycles. We record some of these properties in the following theorem.
Theorem 1.6 ([11, §9, Chapter VI]).
Let $X$ be a smooth variety over $k$. Under the above assumptions, the following hold:
  • (1)
    For any finite étale morphism $f\colon Y\to X$, the étale cycle class map is compatible with the pullback of algebraic cycles along $f$.
  • (2)
    The étale cycle class map is compatible with exterior products of algebraic cycles via the Künneth map in cohomology.
  • (3)
    For two algebraic cycles that intersect transversely, the étale cycle class map is compatible with the intersection product via the cup product in cohomology.
Remark 1.7.
In Theorem 1.6(1), the hypothesis can in fact be weakened to the condition that the codimension of the inverse image of any subvariety of $X$ under a morphism $f\colon Y\to X$ is lower semi-continuous on the fibers. See [5].
2. Motivic cycle class map
2.1. Comparison map
Fix a field $k$. Let $\textbf{Sm}_{k}$ denote the category of smooth schemes over $k$. For each integer $q$, there is a complex (the Voevodsky–Suslin complex) of Nisnevich sheaves with transfers, denoted $\mathbb{Z}(q)$, on $\textbf{Sm}_{k}$ (equipped with the Nisnevich topology). The motivic cohomology of a smooth scheme $X$ is defined through the hypercohomology $H^{p,q}(X,\mathbb{Z}):=\mathbb{H}^{p}_{\mathrm{Nis}}\!\bigl(X,\mathbb{Z}(q)\bigr)$. (Indeed, $\mathbb{Z}(q)$ is also a complex of Zariski sheaves. For a detailed discussion, see [10, Part 3])
Let $\pi\colon X_{\acute{e}t}\to X_{\mathrm{Nis}}$ be the canonical morphism of sites from the étale site to the Nisnevich site. Using the unit of adjunction $\mathrm{id}\longrightarrow R\pi_{*}\pi^{*}$, we obtain a canonical morphism
(2.1) $\mathbb{Z}(q)\longrightarrow R\pi_{*}\pi^{*}\mathbb{Z}(q).$
Since $\pi^{*}\mathbb{Z}(q)$ is the étale sheafification of $\mathbb{Z}(q)$, passing to hypercohomology yields the comparison map $\mathbb{H}^{p}_{\mathrm{Nis}}\!\bigl(X,\mathbb{Z}(q)\bigr)\longrightarrow\mathbb{H}^{p}_{\acute{e}t}\!\bigl(X,\mathbb{Z}(q)\bigr)$. The left-hand side is motivic cohomology, while the right-hand side is known as Lichtenbaum cohomology, which we denote by $H_{L}^{p,q}(X,\mathbb{Z})$. Thus we obtain the comparison map $H^{p,q}(X,\mathbb{Z})\longrightarrow H_{L}^{p,q}(X,\mathbb{Z})$.
For any abelian group $A$, define $A(q):=\mathbb{Z}(q)\otimes_{\mathbb{Z}}A$. This gives motivic and Lichtenbaum cohomology with coefficients in $A$, and hence a comparison map in general:
$H^{p,q}(X,A)\longrightarrow H_{L}^{p,q}(X,A).$
Now consider the case for the group $\Lambda:=\mathbb{Z}/\ell\mathbb{Z}$ with $\ell\neq\mathrm{char}(k)$. There is a quasi-isomorphism in the category of bounded above complexes of étale sheaves
$\mathrm{Cone}\bigl(\mathbb{Z}(1)_{\acute{e}t}\xrightarrow{\ \times\ell\ }\mathbb{Z}(1)_{\acute{e}t}\bigr)\simeq\mu_{\ell}[0],$
[…]
$H_{L}^{p,q}(X,\mathbb{Z}/\ell\mathbb{Z})\cong H^{p}_{\acute{e}t}\!\left(X,\underline{\Lambda}(q)\right).$
where $\underline{\Lambda}(q)$ denotes the $q$th Tate twist for $\underline{\Lambda}:=\underline{\mathbb{Z}/l\mathbb{Z}}$ (See Definition 1.3)
When $p=2q$, the comparison theorem for motivic cohomology in weight $q$ implies that $H^{2q,q}(X,A)\cong\mathop{\mathrm{CH}}\nolimits^{q}(X)\otimes A$ for an abelian group $A$. Consequently, we obtain the motivic cycle class map
(2.2) $cl_{\mathrm{mot}}^{q}\colon\mathop{\mathrm{CH}}\nolimits^{q}(X)\otimes\Lambda\longrightarrow H^{2q}_{\acute{e}t}\!\left(X,\underline{\Lambda}(q)\right).$
where $\Lambda=\mathbb{Z}/l\mathbb{Z}$ with $l\neq\mathop{\mathrm{char}}\nolimits(k)$.
A natural question is the following.
Question 2.3.
Let $X$ be a smooth variety over $k$. Is the motivic cycle class map $cl_{\mathrm{mot}}^{q}$ induced by the étale cycle class map considered in §1.3?
We give a positive answer to this question by carefully examining the construction of the comparison morphism from motivic cohomology to Lichtenbaum cohomology due to Geisser and Levine [3]. Although this result has been known to experts, a detailed reference is hard to find, so we include a proof following their method.
2.2. Naturality of the motivic cycle class map
Let $X$ be a smooth scheme over $k$, and let $\Lambda:=\mathbb{Z}/\ell\mathbb{Z}$ with $\ell\neq\mathrm{char}(k)$. For any integer $q\geq 0$, let $z^{q}(X,*)$ denote Bloch’s cycle complex (of Zariski sheaves) computing higher Chow groups [10, Part 5]. It is known that, after a suitable reindexing, the complex $z^{q}(X,*)$ is quasi-isomorphic to the Voevodsky–Suslin complex $\mathbb{Z}(q)$ (see [3, Section 2.5]). We denote the reindexed complex by $Z^{q}(X,*)$. In particular, its degree-$2q$ term satisfies $Z^{q}(X,2q)=Z^{q}(X)$ , the group of codimension-$q$ algebraic cycles on $X$.
Let $\pi:X_{\text{ét}}\to X_{Zar}$ be the canonical morphism from the étale site to the Zariski site. Then the construction of Geisser and Levine yields a natural map ([3, Equation 3.21])
(2.4) $Z^{q}(X,*)\otimes\Lambda\to R\pi_{*}\pi^{*}(\underline{\Lambda}(q))$
Proposition 2.5.
[3, cf. Theorem 1.1] Under the same assumptions as above, the map on hypercohomology induced by (2.4) coincides with the motivic cycle class map considered in §2.1 (i.e., (2.2)).
Proof.
Since $Z^{q}(X,*)$ is quasi-isomorphic to $\mathbb{Z}(q)$ as complexes of Zariski sheaves, and $\pi^{*}\mathbb{Z}(q)\otimes^{\mathbf{L}}\underline{\Lambda}$ is quasi-isomorphic to $\pi^{*}(\underline{\Lambda}(q))$ as complexes of étale sheaves, we have that the map (2.4) is induced by the unit of adjunction
$\mathrm{id}\longrightarrow R\pi_{*}\pi^{*}$
in the derived category of sheaves. Comparing with (2.1), the uniqueness of the unit of adjunction shows that, after passing to hypercohomology, the map
$H^{p,q}(X,\Lambda)\longrightarrow H^{p,q}_{L}(X,\Lambda)=H^{p}_{\acute{e}t}\!\left(X,\Lambda(q)\right)$
induced by (2.4) coincides with the motivic cycle class map considered in §2.1 ∎
2.3. Motivic versus étale cycle class map
We conclude this section by proving the following theorem. We remind the reader that the proof follows directly from the construction of the motivic cycle class map given by Geisser and Levine [3].
Theorem 2.6.
Let $X$ be a smooth variety over $k$, and let $\Lambda:=\mathbb{Z}/\ell\mathbb{Z}$ with $\ell\neq\mathrm{char}(k)$. Then the motivic cycle class map (2.2) is obtained from the étale cycle class map by tensoring with $\Lambda$ and passing to the quotient by rational equivalence relation.
Proof.
To prove the theorem, it suffices to verify the statement on each graded component of the Chow group of $X$. Recall that $Z^{q}(X,*)$ is the reindexed Bloch cycle complex, with $Z^{q}(X)\otimes\Lambda=Z^{q}(X,2q)\otimes\Lambda$. By Proposition 2.5, we reduce to determining the image of the natural map (cf. (2.4))
(2.7) $\eta:Z^{q}(X,*)\otimes\Lambda\to R\pi_{*}\pi^{*}(\underline{\Lambda}(q))$
As a complex of étale sheaves, $\pi^{*}\!\bigl(\underline{\Lambda}(q)\bigr)$ is quasi-isomorphic to $G^{*}\!\bigl(X,\underline{\Lambda}(q)\bigr)$, the Godement resolution of $\underline{\Lambda}(q)$ ([3, Example 3.3]). Hence the map $\eta$ is reduced to
(2.8) $Z(X,*)\otimes\Lambda\to G^{*}(X,\underline{\Lambda}(q)).$
(See [3, §3.7]) which yields a map $cl^{q}:Z^{q}(X)\otimes\Lambda\to H^{2q}(G^{*}(X,\underline{\Lambda}(q)))$.
For any closed subvariety $W\subset X$ of codimension $q$, let $Z^{q}_{W}(X)$ denote the group of codimension-$q$ cycles on $X$ supported in $W$, and let $G^{*}_{W}\!\bigl(X,\underline{\Lambda}(q)\bigr)$ denote the Godement resolution of $\underline{\Lambda}(q)$ with support in $W$. Clearly, $W$ defines an element of $Z^{q}_{W}(X)$, and we have a natural inclusion $Z^{q}_{W}(X)\subset Z^{q}(X)$.
From [3, Lemma 3.8], we obtain the following commutative diagram:
[…]
where the vertical arrows are the natural maps.
Viewing $W$ as a cycle in both $Z^{q}_{W}(X)$ and $Z^{q}(X)$, we conclude that
$cl^{q}(W)=\bigl(cl^{q}_{X}\otimes\Lambda\bigr)(W),$
which completes the proof.
3. Weil restriction and the motivic cycle class map
3.1. Weil restriction on $\ell$-adic cohomology
In this section, we construct the Weil restriction on $\ell$-adic cohomology. Unless otherwise specified, all cohomology groups will refer to étale cohomology.
We begin by recalling some basic notions. Fix a field $k$ and let $X$ be a scheme over $k$. Let $\ell$ be a prime number with $\ell\neq\mathrm{char}(k)$, and set
$\Lambda_{n}:=\mathbb{Z}/\ell^{n}\mathbb{Z}.$
The $\ell$-adic cohomology of $X$ with coefficients in $\mathbb{Q}_{\ell}$ is defined by
[…]
where the inverse limit is taken with respect to the transition maps
$\underline{\Lambda_{n+1}}(r)\longrightarrow\underline{\Lambda_{n}}(r)$
induced by the natural map $\mathbb{Z}/l^{n+1}\mathbb{Z}\to\mathbb{Z}/l^{n}\mathbb{Z}$.
Notation 3.1.
To simplify notation, we will refer to $\ell$-adic cohomology with coefficients in $\mathbb{Q}_{\ell}$ simply as $\ell$-adic cohomology. Also, we will always assume that $\ell\neq\mathop{\mathrm{char}}\nolimits k$. This will be used in the rest of the paper.
There are two reasons to consider $\ell$-adic cohomology. First, the étale cycle class maps (1.4) and (1.5) do not, in general, extend to Chow groups without additional hypotheses on the dimension of the ambient scheme. In contrast, for $\ell$-adic cohomology, such an extension is always possible [11, Remark 10.8, §10, Chapter VI]. Second, $\ell$-adic cohomology has good descent properties, which are described in the following lemma.
Lemma 3.2.
Let $L/k$ be a finite Galois extension with Galois group $G$, and let $X$ be a scheme over $k$. Then, for all integers $i$ and $r$, there is a canonical isomorphism
$H^{i}\!\left(X_{L},\mathbb{Q}_{\ell}(r)\right)^{G}\;\cong\;H^{i}\!\left(X,\mathbb{Q}_{\ell}(r)\right).$
Proof.
Note that the projection morphism $X_{L}\to X$ is a finite Galois covering morphism (In the sense of [11, Remark 5.4, §5, Chapter I]). Hence we have the Hoschild-Serre spectral sequence
$E_{2}^{p,q}=H^{p}(G,H^{q}(X_{L},\mathbb{Q}_{l}(r)))\Rightarrow H^{p+q}(X,\mathbb{Q}_{l}(r)).$
([11, Theorem 2.20,§2, Chapter III]) Since $H^{q}(X_{L},\mathbb{Q}_{l}(r))$ is a vector space over $\mathbb{Q}_{l}$, which is uniquely $n$-divisible, we have that $E^{p,q}_{2}=0$ except for the term $p=0$ ([4, Corollary 3.3.9, Chapter III]). Hence we have
$H^{q}(X,\mathbb{Q}_{l}(r))\cong H^{0}(G,H^{q}(X_{L},\mathbb{Q}_{l}(r)))=H^{q}(X_{L},\mathbb{Q}_{l}(r))^{G}$
This finishes our proof. ∎
We now turn to the construction of the Weil restriction in $\ell$-adic cohomology. Let $L/k$ be a finite Galois extension with Galois group $G$, and let $X$ be a smooth quasi-projective variety over $L$. We assume that $[L:k]=n$. Still, we denote by $\mathcal{R}(X)$ the Weil restriction of $X$ along $L/k$.
For any cohomology class $f\in H^{i}\!\left(X,\mathbb{Q}_{\ell}(r)\right)$ and any $\sigma\in G$, we denote by $f^{\sigma}$ the corresponding class in
$H^{i}\!\left(X^{\sigma},\mathbb{Q}_{\ell}(r)\right),$
where $X^{\sigma}$ is the conjugate scheme as defined in § 1.1.
Since $X$ is quasi-projective, one can describe $f^{\sigma}$ explicitly using Čech cohomology (see [11, Theorem 2.17, §2, Chapter III]). Let $U\to X$ be an étale covering morphism, and set
$U^{[i+1]}:=\underbrace{U\times_{X}U\times_{X}\cdots\times_{X}U}_{(i+1)\text{ times}}.$
Then a class $f\in H^{i}\!\left(X,\mathbb{Q}_{\ell}(r)\right)$ can be represented by a Čech $i$-cocycle, that is, by a section
$f_{i}\in\Gamma\!\left(U^{[i+1]},\mathbb{Q}_{\ell}(r)\right)$
satisfying the usual Čech cocycle condition. The section $f_{i}$ should be viewed as an étale-locally constant section with values in the sheaf $\mathbb{Q}_{\ell}(r)$. Then $f_{i}^{\sigma}$ corresponds to the same Čech cocycle, but with its values in $\mathbb{Q}_{\ell}(r)$ twisted by the natural Galois action of $\sigma$ on the Tate twist.
Consider the following Künneth product homomorphism ([11, §8, Chapter VI]):
(3.3) $\varphi:\prod_{\sigma\in G}H^{i}(X^{\sigma},\mathbb{Q}_{l}(r))\to H^{ni}(\prod_{\sigma\in G}X^{\sigma},\mathbb{Q}_{l}(nr))$
Now let $f\in H^{i}\!\left(X,\mathbb{Q}_{\ell}(r)\right)$ be a cohomology class. Define
$f^{\prime}:=\varphi\!\left(\prod_{\sigma\in G}f^{\sigma}\right)\in H^{ni}\!\left(\prod_{\sigma\in G}X^{\sigma},\mathbb{Q}_{\ell}(nr)\right),$
Using Čech representatives, one checks that $f^{\prime}$ is $G$-invariant. Since
$\mathcal{R}(X)_{L}\cong\prod_{\sigma\in G}X^{\sigma},$
Lemma 3.2 implies that $f^{\prime}$ descends to a unique class in
$H^{ni}\!\left(\mathcal{R}(X),\mathbb{Q}_{\ell}(nr)\right),$
which we denote by $\mathcal{R}(f)$.
This construction defines a map
$\mathcal{R}\colon H^{i}\!\left(X,\mathbb{Q}_{\ell}(r)\right)\longrightarrow H^{ni}\!\left(\mathcal{R}(X),\mathbb{Q}_{\ell}(nr)\right),$
called the Weil restriction on $\ell$-adic cohomology.
3.2. Compatibility
In this section, we prove the compatibility between the Weil restriction on Chow groups and that on $\ell$-adic cohomology, formulated via motivic cohomology and the motivic cycle class map. Note that the étale cycle class map (1.4) (also for (1.5)) is compatible with the passage to $\ell$-adic cohomology. Hence they extend to maps
(3.4) $cl^{q}_{-}\colon Z^{q}(-)\longrightarrow H^{2q}\!\left(-,\mathbb{Q}_{\ell}(q)\right)$
for any smooth $k$-variety. Moreover, the map (3.4) is compatible with rational equivalence. Hence it descends to a homomorphism
$cl_{-}^{{}^{\prime}\,q}\colon\mathop{\mathrm{CH}}\nolimits^{q}(-)\longrightarrow H^{2q}\!\left(-,\mathbb{Q}_{\ell}(q)\right).$
Note that $cl_{-}^{q}$ and $cl_{-}^{{}^{\prime}\,q}$ have the same image, since an algebraic cycle $\alpha$ and its rational equivalence class $\overline{\alpha}$ map to the same cohomology class. ([11, §10, Chapter VI])
Notation 3.5.
we will, by abuse of notation, use $cl^{q}_{-}$ to denote the $q$th étale cycle class map defined on both algebraic cycles and Chow groups.
Our main result in this section is the following theorem.
Theorem 3.6.
Let $L/k$ be a finite Galois extension of degree $n$ and $X$ be a smooth quasi-projective variety over $L$. For any integer $q$ with $0\leq q\leq\dim X$, the following diagram is commutative:
[…]
where the left vertical map is induced by the Weil restriction on Chow groups, and the right vertical map is the Weil restriction on $\ell$-adic cohomology via the standard identifications in motivic cohomologies.
Under the assumptions of Theorem 3.6, we may apply Theorem 2.6 together with the identifications
$H^{2*,*}(X,\mathbb{Q}_{\ell})\cong\mathop{\mathrm{CH}}\nolimits^{*}(X)\otimes\mathbb{Q}_{\ell}\quad\text{and}\quad H^{2*,*}_{L}(X,\mathbb{Q}_{\ell})\cong H^{2*}\!\left(X,\mathbb{Q}_{\ell}(*)\right)$
to reduce the question to the following proposition.
Proposition 3.7.
Under the assumptions of Theorem 3.6, for any $q\in[0,dim(X)]$, we have the following commutative diagram
[…]
Proof.
Let $G=\mathop{\mathrm{Gal}}(L/k)$. Using Lemma 3.2 and [7, Lemma 2.1], we have that $Z^{nq}(X)\cong Z^{nq}(\prod_{\sigma}X^{\sigma})^{G}$ and $H^{2nq}(\mathcal{R}(X),\mathbb{Q}_{l}(nq))=H^{2nq}(\prod_{\sigma\in G}X^{\sigma},\mathbb{Q}_{l}(nq))^{G}$. Moreover, from the construction of the Weil restriction map, we have that $\mathcal{R}:Z^{q}(X)\to Z^{nq}(\mathcal{R}(X))$ factors through the map
$\varphi:Z^{q}(X)\to Z^{nq}(\prod_{\sigma\in G}X^{\sigma});\alpha\mapsto\prod_{\sigma\in G}\alpha^{\sigma}$
while the map $\mathcal{R}:H^{2q}(X,\mathbb{Q}_{l}(q))\to H^{2nq}(\mathcal{R}(X),\mathbb{Q}_{l}(nq))$ factors through the map
$\psi:H^{2q}(X,\mathbb{Q}_{l}(q))\to H^{2nq}(\prod_{\sigma\in G}X^{\sigma},\mathbb{Q}(nq));f\mapsto\prod_{\sigma}f^{\sigma}.$
As a result, to proves the commutativity of the diagram in the proposition, it suffices to prove the the commutativity of the following diagram
[…]
Now for an algebraic cycle $\alpha\in Z^{q}(X)$, the upper part of the diagram 3.8 gives a cycle $\prod_{\sigma\in G}(cl_{X}^{q}(\alpha))^{\sigma}$ while the lower part gives a cycle $cl_{\prod X^{\sigma}}^{nq}(\prod_{\sigma\in G}\alpha^{\sigma})$. Since the étale cycle class map is compatible with exterior product, we have that
$cl_{\prod X^{\sigma}}^{nq}(\prod_{\sigma\in G}\alpha^{\sigma})=\prod_{\sigma}cl_{X^{\sigma}}(\alpha^{\sigma}).$
Now the statement follows from the compatibility of étale cycle class map with flat pullback (Theorem 1.6) applied to the isomorphism $\sigma:X\to X^{\sigma}$. ∎